a)     Find the moment coefficient of Skewness and kurtosis for the data below.

           9, 3, 4, 2, 9, 5, 8, 4, 7, 4      



1
Expert's answer
2021-11-10T14:26:24-0500

Skewness:


"\mu_3=\frac{\sum (x_i-\mu)^3}{(n-1)\sigma^3}"


"\mu=\sum x_i/n=5.5"


"\sigma=\sqrt{\frac{(x_i-\mu)^2}{n}}=2.42"


"\mu_3=0.302"


kurtosis:


"kurt(X)=\frac{\mu^4}{\sigma^4}"


"\mu^4=\frac{\sum(x_i-\mu)^4}{n}"


"kurt(X)=1.443"